OpenStudy (anonymous):

You have done 3,404 J of work with a force of 37 N on a 46-kg object. Over what distance did you do the work? A. 92 m B. 150 m C. 3131 m D. 1702 m ^_^

OpenStudy (anonymous):

Work= Force* Distance 37 * D=3404J

OpenStudy (anonymous):

3404=37*d so d=3404/37=92m

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

good job